OpenStudy (anonymous):

Find two numberss whose sum is 13 and whose squares total 85?

OpenStudy (lalaly):

x+y=13 x^2 +y^2=85

OpenStudy (lalaly):

x=y-13 substitute it in the second equation to find y

OpenStudy (anonymous):

a+b=13 a^2+b^2=85 2ab=169-85 ab = 84/2=42 a=6 b=7